OpenStudy (anonymous):

An archer releases an arrow from a shoulder height of 1.39 m. When the arrow hits the target 18 m away, it hits point A. When the target is removed, the arrow lands 45 m away. Find the maximum height of the arrow along its parabolic path.

OpenStudy (anonymous):

Ok, we have these points in the parabola: (0, 1.39) and (45, 0)

OpenStudy (anonymous):

but we have no information about initial velocity, angle or the height of the target

OpenStudy (anonymous):

We need to know at least initial velocity or the height of the target in order to obtain another point in the parabola, and, with three points, we can find the parabola equation and find its maximum height.

OpenStudy (anonymous):

i don't see how to solve that, this question does not give any other information or a picture?

OpenStudy (anonymous):

that's a link a the whole thing. I'm on activity 2.

OpenStudy (anonymous):

Ohh now we have a picture!

OpenStudy (anonymous):

So do you think you can help me?

OpenStudy (anonymous):

sorry, i don't see how to find the height of the target...

OpenStudy (anonymous):

how u did the Activity 1?

OpenStudy (anonymous):

how u found the parabola in Activity 1?

OpenStudy (mathmate):

" hits the target at a point 5 ft above the ground" @m4thM1nd @emily0824

OpenStudy (mathmate):



OpenStudy (mathmate):

@emily0824 You still there?

OpenStudy (anonymous):

yes, I'm confused tho

OpenStudy (mathmate):

Can you explain what's confusing you?

OpenStudy (anonymous):

I'm still not sure how to find the maxium height

OpenStudy (mathmate):

You need to find the equation of the trajectory before you can find the maximum height. Do you agree with the 5' height of the target? The information is given in a separate section, but it seems that it is the only information available.

OpenStudy (anonymous):

okay! thanks!

OpenStudy (mathmate):

Are you able to find the trajectory?

OpenStudy (mathmate):

You have three points, and the standard equation of the trajectory is \(f(x)=a(x-h)^2+k\) is this familiar to you?

OpenStudy (mathmate):

with three unknowns to be found: a, h and k.

OpenStudy (mathmate):

We will assume y=0 at ground level.

OpenStudy (anonymous):

yes.. im familiar with that

OpenStudy (mathmate):

Using the three known points (0,1.39), (18, 1.524), and (45, 0) we should be able to solve for the three unknowns a, h and k. ok so far?

OpenStudy (anonymous):

yes

OpenStudy (mathmate):

Do you know the physical interpretation of a, h and k?

OpenStudy (mathmate):

In any case, (h,k) represents the vertex, a is a stretching/compression factor. You would substitute the three points in the equations and solve for a, h and k. 1.39=(0-h)^2+k 1.524=(18-h)^2+k 0=(45-h)^2+k The strategy is to eliminate k from the equations and end up with two equations. Then you solve for a and h. Remember, (h,k) is your vertex. Good luck!

OpenStudy (anonymous):

im not too sure, to be honest i moved schools and this is something i had never done at my old school and this was a chapter i missed. thank you so much for helping!

OpenStudy (mathmate):

Try to work on this and if you need further help, just post what you have. We can check your work. But it is important that you do the work so you don't get stuck on the next problem. The more you work in math, the easier math becomes.